An asset used in a 4-year project falls in the 5-year MACRS class (MACRS Table) for tax purposes. The asset has an acquisition cost of $18,720,000 and will be sold for $4,160,000 at the end of the project.

  

If the tax rate is 23 percent, what is the aftertax salvage value of the asset?
Answer #1
Step-1:Calculation of book value at the end of 4 years
Year Cost Depreciation rate Depreciation expense Accumulated Depreciation expense Book Value
a b c=a*b d e=a-d
1 $ 1,87,20,000.00 20.00% $ 37,44,000.00 $ 37,44,000.00 $ 1,49,76,000.00
2 $ 1,87,20,000.00 32.00% $ 59,90,400.00 $ 97,34,400.00 $     89,85,600.00
3 $ 1,87,20,000.00 19.20% $ 35,94,240.00 $ 1,33,28,640.00 $     53,91,360.00
4 $ 1,87,20,000.00 11.52% $ 21,56,544.00 $ 1,54,85,184.00 $     32,34,816.00
Step-2:Calculation of after tax savage value of the asset
After tax salvage value = Sales Value - ((Sale Value - Book Value)*23%)
= 4160000-((4160000-3234816)*23%)
= $ 39,47,207.68